During a circus act, an elderly performer thrills the crowd by catching a cannonball shot at him. The cannonball has a mass of 10.0 kg and the horizontal component of its velocity is 8.00 m/s when the 65.0-kg performer catches it. If the performer is on nearly frictionless roller skates, what is his recoil velocity?


1
Expert's answer
2021-03-08T08:29:10-0500

Let's use the law of conservation of momentum:


"m_1v_1+m_2v_2=(m_1+m_2)v_{rec},""v_{rec}=\dfrac{m_1v_1+m_2v_2}{(m_1+m_2)},""v_{rec}=\dfrac{10\ kg\cdot8\ \dfrac{m}{s}+0}{(10\ kg+65\ kg)}=1.07\ \dfrac{m}{s}."